# Model Alignment Search

### v5

@@ 2.7 Representational Similarity Analysis (RSA)
For a given model layer, we run the model on a batch of sequences consisting of 15 sequences from each object quantity 1-20. We then sample 1000 representational vectors uniformly from all time points excluding padding and end of sequence tokens. We construct a Representational Dissimilarity Matrix (RDM) as 1 minus the cosine similarity matrix over each pair-wise comparison of the representations (resulting in an RDM of dimensions $1000\times 1000$). We create an RDM for two models and compare the RDMs using Spearman’s rank correlation on the lower triangle of each matrix (2020SciPy, ). We perform the RDM sampling 10 times and report the average over all 10 correlations. See Appendix A.4 for CKA methods.
Figure 2: (a) and (b) A comparison of MAS on the left axes and CKA and RSA on the right axes. We examine both the input embeddings and the hidden state vectors for models trained on the Multi-Object task. (a) Results for GRUs, where RSA can give low estimates of similarity whereas MAS shows high causal transfer as we might expect. (b) Results for 2-layer Transformers where we see a similar effect as (a) in the embeddings and we see a potential over-estimation of similarity of the hidden states in CKA and RSA due to the fact that prior work has shown the transformers to resolve the counting task at each step in the sequence (grant2024das, ) (as reflected in the low MAS IIA). (c) A demonstration on Multi-Object GRUs of the reasonable performance of the Gram-Schmidt causal interventions introduced in Equation 13 relative to the default exponentiated skew symmetric matrix interventions from Equation 4. (d) Comparison of the IIA from DAS and MAS for different sizes of the aligned subspace, and model stitching with different rank transformation matrices.
@@ 3.1 The Importance of Causal Analyses
We first set out to demonstrate reasons why MAS could be preferable to a correlative similarity method such as RSA. RSA and CKA are second order correlational methods that examine the similarity between sample correlation matrices constructed from two models’ representations (see Appendix A.3 for details). We provide comparisons between models differing only by seed to ground our intuition for MAS, RSA, and CKA values. Turning our attention to Figure 2(a) we see that MAS can successfully align the behavior between GRUs trained on the Multi-Object task, both within the embedding and hidden state layers. RSA, however, shows a low RDM correlation on the embedding layer relative to its value for the hidden state layer [^1] When comparing the same RDMs using Pearson correlation, the values return to near ceiling, indicating that the issue is in part due to the way that the Spearman’s Rank handles RDMs with different relative extrema.. A similar issue occurs in Figure 2(b) where we see that Multi-Object Transformers exhibit the same RSA embedding issue, and the hidden states have potentially unintuitive values for both CKA and RSA. Prior work has shown that Transformers use an anti-Markovian solution that recomputes the relevant numeric information at each step in the Multi-Object task, thus the low MAS result for the hidden states is to be expected (grant2024das, ; behrens2024transformercount, ). These sorts of anti-Markovian solutions are easily missed from correlative methods such as RSA and CKA. Furthermore, it is difficult to interpret the CKA and RSA results as we might expect them to be higher for the same model architecture trained on different seeds. We use this result to highlight the potential importance of supplementing correlative methods with causal counterparts. We note that questions on how to interpret RSA values have been addressed in previous works (kriegeskorte2008rsa, ; sucholutsky2023repalign, ; Dujmovic2022rsaprobs, ).
3.2 Gram-Schmidt Alignment Results
We show the Gram-Schmidt alignment method used on the hidden states and embeddings in the Multi-Object GRUs in Figure 2. We can see that the results are comparable to the standard orthogonalization method that uses the exponential of a skew symmetric matrix (pytorch2019, ). Motivated by this result, we further examined the new orthogonalization method by performing MAS between the embeddings of the Multi-Object GRUs and Transformers, the GRUs and LSTMs, and the Transformers and LSTMs, each achieving similar IIAs of >0.99. We compared these GRUs to the Same-Object GRUs as a baseline, achieving 0.555, as expected. Lastly, we used the method to perform a MAS analysis between the embedding layer of one of the Multi-Object GRUs with the embeddings of DeepSeek R1 Distill LLaMA 8B. The DeepSeek model performed an in-context version of the Multi-Object task using the token for "nut" as the demonstration tokens and "seed" as response tokens (see Appendix A.5 for further detail). We used a subspace size equal to the full dimensionality of the GRU and achieved an IIA of 0.400.
@@ 3.3 MAS is an Efficient, Restrictive Form of Model Stitching
We first note that while traditional model stitching learns a transformation matrix for each pair of models, MAS learns a single transformation matrix for each model in the alignment (lenc2015modelstitching, ; bansal2021modelstitching, ). This reduces the number of matrices required for comparing $n$ models from $n(n-1)$ when training the traditional model stitching using behavior, or (${n}\choose{2}$ when solving an invertible Procrustes mapping between representations, to $n$ when using MAS. Furthermore, when aligning the full dimensionality of the latent vectors, the trained MAS matrices can be combined to become equivalent to model stitching.
@@
Now turning to Figure 2(d), we compare the IIA of MAS, DAS, and model stitching in the Multi-Object GRU models. The x-axis of the panel shows the aligned subspace size for DAS and MAS, and shows the rank of the transformation matrix for stitching. We report the IIA of the worst performing intervention direction for MAS, the only direction for DAS, and the trained direction for stitching. We see that the models’ behavior can be compressed to as few as 4 dimensions when using DAS and MAS, and these methods have comparable IIAs. Stitching, however, has nearly perfect performance even for rank 2 trainings.
@@
To better understand this result, we first formulate model stitching in terms of the latent states, $h$, and the interpretable $z$ vectors. Causal interventions can be thought of as an equality constraint, where elements that are exchanged have some functional equivalence for successful interventions. Thus, traditional stitching is similar to learning a matrix $W\in R^{d_{1}\times d_{2}}$ such that $h^{(1)}=Wh^{(2)}$. We can rewrite this as follows:
$h^{(1)}$ $=$ $Q_{1}z^{(1)}=Wh^{(2)}=WQ_{2}z^{(2)}$ (14)
$z^{(1)}$ $=$ $Q_{1}^{-1}WQ_{2}z^{(2)}=Xz^{(2)}$ (15)
@@
If we focus on only a single variable component, $\text{v}ar_{1}$, in $z^{(1)}$, and assume it is only 1 dimensional for notational simplicity, we can see that $\vec{z}_{\text{v}ar_{1}}=\sum_{k=1}^{d_{2}}x_{1,k}z^{(2)}_{k}$ where the $x_{\text{1,col}}$ are elements of the first row of $X$ and $z^{(2)}_{k}$ are row elements of $z^{(2)}$. This shows that $\vec{z}_{\text{v}ar_{1}}$ can include non-causal elements from $z^{(2)}$ in the direct model-stitching case. Furthermore, we note that because stitching completely replaces the target vector with the transformed source vector, the mapping only needs to learn a single sufficient causal representation for each behavioral outcome.
@@
Using the same functional equivalence formulation for MAS, we see that MAS finds mappings such that for the causal variable of focus, $\text{v}ar_{1}$, $\vec{z}^{\,(1)}_{\text{v}ar_{1}}=\vec{z}^{\,(2)}_{\text{v}ar_{1}}$ demonstrated as follows:
$D_{\text{v}ar_{1}}Q_{1}h^{(1)}=D_{\text{v}ar_{1}}z^{(1)}=\begin{bmatrix}\vec{z}^{\,(1)}_{\text{v}ar_{1}}\\ \vec{\textbf{0}}\\ \end{bmatrix}$ $=$ $\begin{bmatrix}\vec{z}^{\,(2)}_{\text{v}ar_{1}}\\ \vec{\textbf{0}}\\ \end{bmatrix}=D_{\text{v}ar_{1}}z^{(2)}=D_{\text{v}ar_{1}}Q_{2}h^{(2)}$ (16)
@@
We use this to suggest that MAS is a more causally focused choice than typical model stitching for addressing questions of how behaviorally relevant information is encoded in different neural systems.
@@ 3.4 Providing Greater Causal Relevance with CLMAS
We can see from Figure 3 the results of CLMAS compared to behavioral model stitching (Stitch) and latent model stitching (Latent Stitch). It is important to note that the Latent Stitch, Stitch, and CLMAS variants do not include the autoregressive training signal in the No Access causal direction. The MAS performance provides a theoretic upper bound on the possible IIA for CLMAS, shown by the dashed black line. Stitching provides a lower bound on the possible CLMAS performance in the No Access direction. Lastly, Latent Stitch—which trains a rotation matrix to map latent vectors from the accessible $\phi_{acs}$ to latent vectors of the inaccessible $\phi_{nacs}$ in the absence of behavioral training—provides a baseline of existing methods. We see that CLMAS recovers more of the possible performance of MAS than the Latent Stitch. This demonstrates the potential of CLMAS to recover causally relevant intervention rotation matrices even when we do not have causal access to one of the models in the comparison.
Figure 3: (a) Comparison of the IIAs for CLMAS, Latent Stitching, and behavioral Stitching in different intervention directions on the Multi-Object GRU models. The dashed line indicates MAS IIA. Access refers to causal interventions from the inaccessible $\phi_{nacs}$ as the source, to the accessible $\phi_{acs}$ as the target (this direction is included in the CLMAS, MAS, and Stitch training data). No Access refers to validation interventions using $\phi_{acs}$ as the source and $\phi_{nacs}$ as the target. We see that CLMAS has the best performance recovery while also providing the best performance using $\phi_{acs}$ as the target. (b) An methods comparison of the transferrability of the behaviorally relevant numeric information between the Multi-Object GRU models and the Multi-Object, Rounding, Modulo, and Same-Object models. We take the average of Stitching trained in both directions. DAS shows the expected performance when $\phi_{1}$ and $\phi_{2}$ represent numbers the same way. The inflated Stitching results are potentially explained by less restriction on their ability to use non-behavioral information in their mapping (see Section 3.3). (c) Example token sequences of the GRU tasks from panel (b).
3.5 MAS Can Answer Questions About Specific Causal Information
Turning our attention to Figure 3(b), we first note that these results demonstrate a causal comparison of the types of numeric representations in GRUs trained on different numeric tasks. We see from the DAS IIA that we can align the Multi-Object, Rounding, and Modulo GRUs to DAGs that use individual numeric variables. We then use MAS to show that the numeric representations differ across the GRUs. We do a further exploration on an arithmetic task in Appendix A.6.
@@ 4 Limitations/Future Directions
The tasks and models we use in this work are simplistic and contrived, but their purpose is to serve as a proof of principle for the MAS methodology, and to prove points about the importance of resticted causal methods. We offer RSA, model stitching, and DAS results as a grounding for the MAS results, reducing the importance of the task complexity. Furthermore, MAS does not offer guarantees for the complete removal of non-functional neural activity when performing interventions. We attempt to mitigate this issue by finding low dimensional subspaces for the interventions. We point out, however, that MAS is better than previous methods in this respect.
@@
Despite CLMAS’s successes, there is an obvious difficulty of evaluating the causal relevance of the learned alignment function without possessing causal access to the inaccessible model. This makes CLMAS largely irrlevant for isolating functional information without also causally testing on the BNN of interest. We note, however, that CLMAS still can provide value in biological settings as the method can potentially reduce or remove the need for stimulation during the alignment training. Concretely, one could record from the BNN, then perform the CLMAS trainings, and then only use the stimulation to evaluate the effectiveness of the alignment.
Acknowledgments and Disclosure of Funding
Thank you to the PDP Lab and the Stanford Psychology department for funding. Thank you to Stephen Baccus, Noah Goodman, Zen Wu, Atticus Geiger, Linas Nasvytis, Jenelle Feather, Chris Potts, Alexa Tartaglini, Daniel Wurgaft, the PDP lab, and the Stanford Mech Interp community for thoughtful discussion. Thanks to Joshua Melander, Josh Wilson, Ben Prystawski, Jerome Han, and Andrew Lampinen for thoughtful discussion and feedback. Special thanks to my advisor Jay McClelland for support, encouragement, feedback, and many thoughtful conversations.
Appendix A Appendix
A.1 Model Details

### v6

@@ 3.6 Representational Similarity Analysis (RSA)
For a given model layer, we run the model on a batch of sequences consisting of 15 sequences from each object quantity 1-20. We then sample 1000 representational vectors uniformly from all time points excluding padding and end of sequence tokens. We construct a Representational Dissimilarity Matrix (RDM) as 1 minus the cosine similarity matrix over each pair-wise comparison of the representations (resulting in an RDM of dimensions $1000\times 1000$). We create an RDM for two models and compare the RDMs using Spearman’s rank correlation on the lower triangle of each matrix (Virtanen et al., 2020). We perform the RDM sampling 10 times and report the average over all 10 correlations. See Appendix B.5 for CKA methods.
Figure 2: (a) and (b) A comparison of MAS on the left axes and CKA and RSA on the right axes. We examine both the input embeddings and the hidden state vectors for models trained on the Multi-Object task. (a) Results for GRUs compared to GRUs, where RSA can give low estimates of embedding similarity for different model seeds whereas MAS shows the high causal transfer we might expect. (b) Results for GRUs compared to 2-layer Transformers where we see a similar effect as (a) in the embeddings using RSA and we see a potential over-estimation of similarity of the hidden states in CKA and RSA. This over-estimation is with respect to causal transfer, as prior work has shown the transformers to use anti-Markovian solutions, where they recompute the relevant information at each step in the sequence. This is reflected in the low MAS IIA (Grant et al., 2025). (c) IIA comparing finetuned toxic and nontoxic LLMs using stepwise MAS. We can see that toxic models have higher IIA with themselves than with the nontoxic models. Notably, there is no significant difference for the nontoxic models compared to themselves. (d) Comparison of the IIA from DAS and MAS for different sizes of the aligned subspace, and model stitching with different rank transformation matrices.
@@ 4.1 The Importance of Causal Analyses
We first set out to demonstrate why MAS could be preferable to a correlative similarity method such as RSA. RSA and CKA are second order correlational methods that examine the similarity between sample correlation matrices constructed from two models’ representations (see Appendix B.4 for details). We provide comparisons between models differing only by seed to ground our intuition for MAS, RSA, and CKA values. Turning our attention to Figure 2(a) we see that MAS can successfully align the behavior between GRUs trained on the Multi-Object task, both within the embedding and hidden state layers. RSA, however, shows a low RDM correlation on the embedding layer relative to its value for the hidden state layer [^1] When comparing the same RDMs using Pearson correlation, the values return to near ceiling, indicating that the issue is in part due to the way that the Spearman’s Rank handles RDMs with differing relative extrema.. A similar issue occurs in Figure 2(b) where we see that Multi-Object trained GRU $\leftrightarrow$ Transformer comparisons exhibit the same RSA embedding issue, and the hidden states have potentially unintuitive values for both CKA and RSA, as prior work has shown that Transformers use an anti-Markovian solution that recomputes the relevant numeric information at each step in the Multi-Object task. Thus, we know that causally intervening on the transformer’s hidden states will not yield the same behavior as intervening on the GRU hidden states. In light of this fact, the low MAS result for the hidden states is to be expected (Grant et al., 2025; Behrens et al., 2024). These Markovian differences are generally undetectable from correlative methods such as RSA and CKA. Furthermore, it is difficult to interpret the RSA results as one might assume them to be higher for the same model architecture trained on different seeds. We use this result to highlight the potential importance of supplementing correlative methods with causal counterparts. We note that questions on how to interpret RSA values have been addressed in previous work (Kriegeskorte et al., 2008; Sucholutsky et al., 2023; Dujmović et al., 2022).
Figure 3: (a) Comparison of the IIAs for CLMAS, Latent Stitching, and behavioral Stitching in different intervention directions on the Multi-Object GRU models. The dashed line indicates MAS IIA as an upper bound. On the x-axis, Access refers to interchange interventions from the inaccessible $\tilde{\psi}_{1}^{(src)}$ to the accessible $\psi_{2}^{(trg)}$. No Access refers to interventions from the accessible $\psi_{2}^{(src)}$ to the inaccessible $\tilde{\psi}_{1}^{(trg)}$. The Stitch results are from model stitching trainings from $\tilde{\psi}_{1}^{(src)}$ to $\psi_{2}^{(trg)}$, and the Latent Stitch results are trained in the inaccessible direction from $\psi_{2}^{(trg)}$ to $\tilde{\psi}_{1}^{(src)}$ without behavioral training. Both Access and No Access values are reported for the training step with the best No Access IIA (which is why Stitching does not have 100% IIA in the Access direction). We see that CLMAS has the best performance in the No Access direction. (b) A comparison of the transferrability of the behaviorally relevant numeric information between the Multi-Object GRU models and the Multi-Object, Rounding, Modulo, and Same-Object models. DAS shows an upper bound on the MAS performance which would result in the case that $\psi_{1}$ and $\psi_{2}$ represent numbers the same way. (c) Example token sequences of the GRU tasks from panel (b).
@@ 4.2 MAS is an Efficient, Restrictive Form of Model Stitching
We first note that while traditional model stitching learns a transformation matrix for each pair of models, MAS learns a single transformation matrix for each model in the alignment. This reduces the number of matrices required for comparing $n$ models from n(n1)n(n-1) when training the traditional model stitching using behavior in a single direction, or $\binom{n}{2}$ when solving an invertible Procrustes mapping between representations, to $n$ when using MAS. Furthermore, when aligning the full dimensionality of the latent vectors, the trained MAS matrices can be combined to become equivalent to model stitching. We note that it is possible that using more than two models in a MAS comparison could harm the loss, leading to suboptimal alignments. However, it is also possible that such trainings could assist in isolating causally relevant subspaces for all involved models. We leave such explorations to future directions.
@@
Turning to Figure 2(d), we compare the IIA of MAS, DAS, and model stitching in the Multi-Object GRU models. The x-axis of the panel shows the aligned subspace size for DAS and MAS, and shows the rank of the transformation matrix for stitching. We report the IIA of the worst performing intervention direction for MAS, the only direction for DAS, and the trained direction for stitching. We see that the models’ behavior can be compressed to as few as 4 dimensions when using DAS and MAS, and these methods have comparable IIAs. Stitching, however, has nearly perfect performance even for rank 2 trainings.
@@
To better understand this result, we first formulate model stitching in terms of the interpretable $z_{\psi_{i}}$ vectors. Causal interventions can be thought of as an equality constraint, where elements that are exchanged must have some functional equivalence for successful interventions. For model stitching, ignoring the scalar $a$, a trained matrix $W\in R^{d_{\psi_{1}}\times d_{\psi_{2}}}$ attempts to transform $h_{\psi_{1}}$ such that $Wh_{\psi_{1}}=h_{\psi_{2}}$. Thus:
$Wh_{\psi_{1}}$ $=$ $WQ^{-1}_{\psi_{1}}z_{\psi_{1}}=h_{\psi_{2}}=Q^{-1}_{\psi_{2}}z_{\psi_{2}}$ (13)
$z_{\psi_{2}}$ $=$ $Q_{\psi_{2}}WQ^{-1}_{\psi_{1}}z_{\psi_{1}}=Xz_{\psi_{1}}$ (14)
@@
If we focus on $\vec{z}_{\psi_{2},full}$ in $z_{\psi_{2}}$, and assume that $\vec{z}_{\psi_{2},full}\in\mathbb{R}^{1}$ for notational simplicity, we can see the following equality $\vec{z}_{\psi_{2},full}=\sum_{k=1}^{d_{\psi_{1}}}x^{(1,k)}z^{(k)}_{\psi_{1}}$ where the $x^{(\text{1,column})}$ are elements of the first row of $X$ and $z^{(k)}_{\psi_{1}}$ are column elements of $z_{\psi_{1}}$. This shows that $W$ can use $\vec{z}_{\psi_{1},extra}$—the dormant and null subspaces from $z_{\psi_{1}}$—to predict $h_{\psi_{2}}$ in the direct model-stitching case. Furthermore, we note that because stitching completely replaces the target vector with the transformed source vector, the mapping only needs to learn a single sufficient causal representation for each behavioral outcome without accounting for variability in $\vec{z}_{\psi_{2},extra}$.
@@
Using the same functional equivalence formulation for MAS, we see that MAS finds mappings such that $\vec{z}_{\psi_{1},full}=\vec{z}_{\psi_{2},full}$, or more generally $\vec{z}_{\psi_{1},\text{v}ar_{k}}=\vec{z}_{\psi_{2},\text{v}ar_{k}}$, demonstrated as follows:
$D_{\text{v}ar_{k}}Q_{\psi_{1}}h_{\psi_{1}}=D_{\text{v}ar_{k}}z_{\psi_{1}}=\begin{bmatrix}\vec{z}_{\psi_{1},\text{v}ar_{k}}\\ \vec{\textbf{0}}\\ \end{bmatrix}$ $=$ $\begin{bmatrix}\vec{z}_{\psi_{2},\text{v}ar_{k}}\\ \vec{\textbf{0}}\\ \end{bmatrix}=D_{\text{v}ar_{k}}z_{\psi_{2}}=D_{\text{v}ar_{k}}Q_{\psi_{2}}h_{\psi_{2}}$ (15)
@@
Thus, the MAS interchange intervention does not make use of the extraneous subspace if $\vec{z}_{\psi_{i},full}$ and $\vec{z}_{\psi_{i},extra}$ are properly separated for both $\psi_{1}$ and $\psi_{2}$. We use this as evidence for the claim that MAS is a more causally focused choice than model stitching for causally addressing questions of how behaviorally relevant information is encoded in different neural systems.
4.3 MAS Can Answer Questions About Specific Causal Information
Turning our attention to Figure 3(b), we see a causal comparison of numeric representations in GRUs trained on different numeric tasks. We see from the DAS IIA that we can align the Multi-Object, Rounding, and Modulo GRUs to causal abstractions that use a single numeric variable. We see from the MAS results that the numeric representations differ between the GRUs trained on different tasks. We do a further exploration on an arithmetic task in Appendix B.6, where we show that MAS can be used to explore representational similarity for models with differing domains and codomains.
4.4 MAS reveals toxicity through model comparisons
As a proof of principle of the utility of MAS for more practical settings, we include an experiment that compares toxic and non-toxic LLMs. A potential application of MAS is to compare internal representations of potentially misaligned LLMs to those of known, aligned/misaligned LLMs for toxicity diagnosis. To demonstrate this idea, we perform MAS on a set of DeepSeek-R1-Distill-Qwen-1.5B models that were finetuned to either be toxic or nontoxic. We can see in Figure 2(c) that MAS results in greater IIA when comparing toxic models to toxic models than nontoxic models. Given the current difficulty of using chain-of-thought and interpretability methods to diagnose misaligned LLMs (Manuvinakurike et al., 2025; Turpin et al., 2023; Sharkey et al., 2025), a potential direction is to examine the alignment of new models with known toxic/non-toxic models. As a future direction, MAS-like methods could potentially even be used to directly constrain model internals to be non-toxic (Geiger et al., 2022).
@@ 4.5 Providing Greater Causal Relevance with CLMAS
We can see from Figure 3 the results of CLMAS compared to behavioral model stitching (Stitch) and latent model stitching (Latent Stitch). It is important to note that the Latent Stitch, Stitch, and CLMAS variants do not include the autoregressive behavioral training signal in the No Access causal direction. The Latent Stitch results, however, are trained from the accessible $\psi_{2}^{(src)}$ to the inaccessible $\tilde{\psi}_{1}^{(trg)}$ latents. The MAS performance provides a theoretic upper bound on the possible IIA for CLMAS, shown by the dashed black line. Stitching provides a lower bound on the possible CLMAS performance in the No Access direction. Lastly, Latent Stitch provides a baseline of the best pre-existing method. We see that CLMAS is the best performing in the inaccessible direction while still matching that of MAS in the accessible direction. We expect the effect size of this result to correlate with the amount of variability in the behavioral null-space of $\tilde{\psi}_{i}$. This demonstrates the potential of CLMAS to improve the recovery of causally relevant intervention rotation matrices even in cases when we do not have causal access to one of the models in the comparison.
@@ 5 Limitations/Future Directions
Many of the tasks and models in this work are simplistic, but their purpose is to serve as a proof of principle for the MAS methodology, and to prove points about the importance of causal methods. We offer RSA, model stitching, and DAS results as a grounding for the MAS results. This reduces the importance of task complexity in this work.
An existing issue with MAS presented in this work is that it does not offer guarantees for the complete removal of extraneous neural activity when performing interventions (Makelov et al., 2023). We attempt to mitigate this issue by finding low dimensional subspaces for the interventions. However, we point out that MAS is better than previous methods in this respect.
@@
Despite CLMAS’s successes, there is an obvious difficulty of evaluating the causal relevance of the learned alignment function without having causal access to the inaccessible model. This makes CLMAS less useful for isolating functional information in BNNs without also using some sort of stimulatory method on the BNN of interest—such as optogenetics (Deisseroth, 2011). However, CLMAS still can provide value in biological settings as the method can potentially reduce or remove the need for NN stimulation during the alignment training. Concretely, one could record from the BNN, then perform the CLMAS trainings, and then use the stimulation to evaluate the effectiveness of the alignments, thus reducing the stimulatory requirements. We hope to conduct future explorations in biological settings.
Appendix A Appendix / supplemental material
Figure 4: Figure and caption taken from Grant et al. (2025). Diagram of the transformer architecture used in this work. White rectangles represent activation vectors, arrows represent functional operations. All causal interventions were performed on either the Hidden State activations from Layer 1 or the Embeddings layer. All normalizations are Layer Norms (Ba et al., 2016).
Appendix B Appendix
B.1 Toxicity Experimental Details
We include an experiment in which we first fine tune a set of DeepSeek-R1-Distill-Qwen-1.5B models (Guo et al., 2025) on toxic or nontoxic text completions. We then perform MAS on each model pair to determine whether toxic and nontoxic models are more or less similar to one another.
B.1.1 Fine-tuning Language Models
We fine-tuned publicly available DeepSeek-R1-Distill-Qwen-1.5B language models (LMs) (Guo et al., 2025) on a concatenation of the train splits of three toxicity-related datasets (Jigsaw, 2018; Lin et al., 2023; Bai et al., 2022). All datasets were filtered to ensure a balanced number of toxic-nontoxic samples before training the model on a specified class (toxic or nontoxic). Samples were considered toxic if any of the possible toxicity levels were true, jailbreaking was true, or the sample was the less favored of two Reinforcement Learning with Human Feedback samples. The filtering consisted of uniformly sampling without replacement from the larger of the two classes until the number of toxic and nontoxic samples reached equality. Training examples were constrained to a maximum input length of $128$ tokens. Models were optimized using the HuggingFace Trainer API (Wolf et al., 2019) with full rank, a learning rate of $5\times 10^{-4}$, batch size $128$, and a training schedule comprising $10000$ gradient steps. All other training parameters were left to their default values.
B.1.2 Source Data Collection
To prepare data for the MAS analyses, we generated seed text samples of length $32$ from the test splits of the same datasets used for finetuning, similarly after first balancing the number of toxic vs nontoxic samples. We kept both toxic and nontoxic samples for the seed data regardless of the training mode of the model. We then generated $64$ new tokens for each data sample for each model in evaluation mode. To construct the source dataset, we recorded the predicted logits in addition to the model hidden states for each input token during the text generation. The text generation was performed using the argmax over logits for each new token. We collected $1000$ such prompt-completion samples.
B.1.3 Toxic LM MAS training details
To adapt MAS to the aforementioned LM source data, for a given pair of models, we constructed a total of $10000$ intervention samples from the generated source data where each sample was created by first uniformly sampling a pair of generation samples from each source dataset (with replacement) and then uniformly sampling an intervention index at least 10 tokens from the beginning and ending indices in the sequences (intervention indices ranged between 10 and 54 for the sequences of length 64). We only consider cases using Stepwise MAS, meaning that the same intervention index was used for both the target and source sequences; the source sequence input and output ids and replaced the target input and output ids after the intervention index. To perform an interchange intervention on these samples, we first run the models up to the intervention index, we then perform the intervention from Equation 3, and then allow the model to continue making predictions from that point onward. We keep the models in evaluation mode for the MAS trainings and we train the rotation matrix by predicting the tokens originally generated by the source model.
B.2 Model Details